40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6133     Entry time: Sun Dec 18 18:45:22 2011
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Y arm ALS : time series and noise budget 
As I said in the previous entry (#6126) my current goals were :
 (1) Take a noise budget when the standard ALS configuration is applied
 (2) Take a beautiful time series to show how ALS brings the cavity to the resonance point

Here are the latest plots that I have obtained from the Friday night:

    Time Series   

time_series.png

 The data starts from a point where the cavity is kept away from the resonance point by 200 kHz (in terms of the green laser's frequency).
Then 30 sec after, a cavity sweep started until the main laser becomes resonant for the arm cavity.
After 2.5 minutes the sweep was quit and the arm length was held at this point to show the
stability of the ALS servo.
 
         Noise Budget         

Yarm_ALS_2011DEC17.png

The residual motion in the arm displacements reached 70 pm in rms.

Note that the UGF was at about 100 Hz.
One of the improvements we made in the Friday was the removal of the 60 Hz line noise (#6127).
Currently the rms is dominated by two components:
     (1) A bump around 10 Hz, which is due to lack of the servo gain around there.
         => This can be improved by optimizing the servo filter shape
     (2) High frequency noise above 40 Hz.
         => This can be improved by either decreasing the noise itself or lowering the UGF.
ELOG V3.1.3-